# onlineReferences

### v5

@@ 2   DID-IV in two time periods
In this section, we formalize an instrumented difference-in-differences (DID-IV) in two-period/two-group settings. We first establish the target parameter and the identifying assumptions in this design. We then investigate the treatment adoption behavior across units, and clarify the interpretation of the parallel assumption in the outcome. Finally, we describe the connections between DID-IV and Fuzzy DID proposed by de Chaisemartin and D’Haultfœuille (2018).
@@ 2.1   Set up
In practice, researchers are interested in the effect of a treatment $D_{i,t}$ on an outcome $Y_{i,t}$, and the instrument $Z_{i,1}$ typically represents a program or policy shock that encourages people to adopt the treatment in period $t=1$. For instance, Duflo (2001) estimates returns to schooling in Indonesia, exploiting variation arising from a new school construction program across regions and cohorts as an instrument for education attainment.
Next, we introduce the potential outcomes framework. Let $Y_{i,t}(d,z)$ denote the potential outcome in period $t$ when unit $i$ receives the treatment path $d\in\mathcal{S}(D)$ and the instrument path $z\in\mathcal{S}(Z)$. Similarly, let $D_{i,t}(z)$ denote the potential treatment status in period $t$ when unit $i$ receives the instrument path $z\in\mathcal{S}(Z)$. We refer to $D_{i,t}((0,0))$ as unexposed treatment and $D_{i,t}((0,1))$ as exposed treatment. Since the treatment and the instrument takes only two values, one can write the observed treatments $D_{i,t}$ and outcomes $Y_{i,t}$ as follows.
@@
Assumption 1 states that potential outcomes $Y_{i,t}(d,z)$ depend only on the current treatment status $d_{t}$ and the instrument path $z$. In the DID literature, several studies adopt this assumption in non-staggered treatment settings (e.g. de Chaisemartin and D’Haultfœuille (2020); Imai and Kim (2021)). [^4] When the treatment is binary and only two periods exist, one can explicitly introduce the treatment path in potential outcomes, as it takes only four values. We therefore consider this case in online Appendix Section A.3 as an extension. One can also weaken Assumption 1 in other cases, such as non-binary, ordered treatment, or multiple time periods, but it requires cumbersome notations and complicates the definition of the target parameter, being not practical for empirical applications.
@@ 2.2   Target parameter in $2\times 2$ DID-IV designs
This parameter measures the treatment effects in period $1$, for those who belong to an exposed group ($E_{i}=1$), and are induced to treatment by instrument in period $1$. In the DID-IV literature, de Chaisemartin (2010) considers the same target parameter, and Hudson et al. (2017) define the local treatment effect (LATE) in period $t=1$ as their target parameter, which is unconditional on $E_{i}$. The LATET has been also proposed in heterogeneous effects IV models with binary instrument (e.g. Słoczyński (2020); Słoczyński et al. (2022)).
We define the LATET as the target parameter for two reasons. First, this parameter would be particularly of interest if the instrument reflects a policy change of interest to the researcher (Heckman and Vytlacil (2001), Heckman and Vytlacil (2005)). Second, the LATET is a natural extension of the target parameter in DID designs, the so-called average treatment effects on the treated (ATT); both causal parameters measure the treatment effects among the units affected by a policy shock (the treatment in DID designs) and belonging to an exposed group (treatment group).
@@ 2.3   Identification assumptions in $2\times 2$ DID-IV designs
Given Assumption 1 and Assumption 2, we can write the observed outcomes $Y_{i,t}$ as
@@
Assumption 1 and Assumption 2 also allow us to introduce the notion of exposed and unexposed outcomes. For any $z\in\mathcal{S}(Z)$, let $Y_{i,t}(D_{i,t}(z))$ denote the outcome if the instrument path is $z$:
@@
Hereafter, we refer to $Y_{i,t}(D_{i,t}((0,0)))$ as unexposed outcomes and $Y_{i,t}(D_{i,t}((0,1)))$ as exposed outcomes. [^5] Exposed and unexposed outcomes are not new concepts in econometrics. For example, under IV designs, we compare the expectation of the outcome with instrument and the one without instrument in the numerator of the wald estimand: $E[Y|Z=1]-E[Y|Z=0]$. This comparison corresponds to the one between the expectation of the exposed outcome and the expectation of the unexposed outcome; that is, $E[Y|Z=1]-E[Y|Z=0]=E[Y(D(1))|Z=1]-E[Y(D(0))|Z=0]$.
@@
Assumption 6 is a parallel trends assumption in the outcome, requiring that the evolution of the unexposed outcome is, on average, the same between exposed and unexposed groups. For instance, in Duflo (2001), this assumption requires that the expectation of log annual earnings would have followed the same path from period 00 to period $1$ between exposed and unexposed groups in the absence of a policy shock.
Note that Assumption 6 depends on the unexposed outcome, and does not depend on the potential outcomes directly. In Subsection 2.4, however, we show that the average time trends of the unexposed outcome in Assumption 6 can be written as the weighted average of the expectation of the potential outcomes, and this assumption can be interpreted as the restriction on the time effect between exposed and unexposed groups.
When empirical researchers exploit variation arising from a policy shock as an instrument for treatment and use the Wald-DID estimand, they often refer to Assumption 5 and 6. For instance, Duflo (2001) estimates returns to schooling in Indonesia, relying on “the identification assumption that the evolution of wages and education across cohorts would not have varied systematically from one region to another in the absence of the program”.
Finally, we assume a relevance condition. This assumption guarantees that the Wald-DID estimand $w_{DID}$ is well defined. [^8] Note that Assumption 3, 4, 5, and 7 imply that the share of compliers in period $1$ is non-zero. $E[D_{i,1}-D_{i,0}|E_{i}=1]-E[D_{i,1}-D_{i,0}|E_{i}=0]=Pr(CM^{Z}|E_{i}=1)>0.$ This ensures that the LATET in period $t=1$ is also well defined.
@@
2.4   Interpreting the parallel trends assumption in the outcome
In this subsection, we clarify the interpretation of the parallel trends assumption in the outcome in DID-IV designs. It is less clear than that in DID designs because this assumption does not depend on the untreated outcome, but depends on the unexposed outcome.
First, we introduce the additional notation. Let $G_{i}^{T}=(D_{i,0}((0,0)),D_{i,1}((0,0)))$ denote the group variable for unit $i$ that represents the treatment adoption process during the two periods if the instrument path is $z=(0,0)$. Recall that in DID-IV designs, the units are allowed to take the treatment both period 00 and $1$ without instrument. In this setting, one can see that time also plays the role of instrument; that is, time also affects one’s treatment choices during the two periods. Similar to the group variable $G_{i}^{Z}$ introduced in Subsection 2.1, we define $G_{i}^{T}=(0,0)\equiv NT^{T}$ as the time never-takers, $G_{i}^{T}=(1,1)\equiv AT^{T}$ as the time always-takers, $G_{i}^{T}=(0,1)\equiv CM^{T}$ as the time compliers, and $G_{i}^{T}=(1,0)\equiv DF^{T}$ as the time defiers.
Table 1: Exposed group (assigned to $z=(0,1)$)
observed counterfactual
$(D_{0}((0,0)),D_{1}((0,1)))$ $D_{1}((0,0))=1$ $D_{1}((0,0))=0$
$D_{0}((0,0))=1,D_{1}((0,1))=1$ $AT^{Z}\land AT^{T}$ $CM^{Z}\land DF^{T}$
$D_{0}((0,0))=1,D_{1}((0,1))=0$ $DF^{Z}\land AT^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}((0,0))=0,D_{1}((0,1))=1$ $AT^{Z}\land CM^{T}$ $CM^{Z}\land NT^{T}$
$D_{0}((0,0))=0,D_{1}((0,1))=0$ $DF^{Z}\land CM^{T}$ $NT^{Z}\land NT^{T}$
Table 2: Unexposed group (assigned to $z=(0,0)$)
observed counterfactual
$(D_{0}((0,0)),D_{1}((0,0)))$ $D_{1}((0,1))=1$ $D_{1}((0,1))=0$
$D_{0}((0,0))=1,D_{1}((0,0))=1$ $AT^{Z}\land AT^{T}$ $DF^{Z}\land AT^{T}$
$D_{0}((0,0))=1,D_{1}((0,0))=0$ $CM^{Z}\land DF^{T}$ $NT^{Z}\land DF^{T}$
$D_{0}((0,0))=0,D_{1}((0,0))=1$ $AT^{Z}\land CM^{T}$ $DF^{Z}\land CM^{T}$
$D_{0}((0,0))=0,D_{1}((0,0))=0$ $CM^{Z}\land NT^{T}$ $NT^{Z}\land NT^{T}$
Notes: These tables represent mutually exclusive and exhaustive types under DID-IV designs. The types painted in gray color are excluded by Assumption 3 (monotonicity assumption).
By exploiting the group variables $G_{i}^{Z}$ and $G_{i}^{T}$, we can describe how the identifying assumptions in $2\times 2$ DID-IV designs restrict the treatment adoption behavior across units in exposed and unexposed groups, respectively. Table 1 and Table 2 show that we can partition units in exposed and unexposed groups into six mutually exclusive and exhaustive types under $2\times 2$ DID-IV designs. [^9] In an exposed group, we can observe $(D_{0}((0,0)),D_{1}((0,1)))$, but not $D_{1}((0,0))$ for each unit (see Table 1). We use Assumption 4 (no anticipation assumption in the first stage), replacing $D_{0}((0,1))$ with $D_{0}((0,0))$. In an unexposed group, we can observe $(D_{0}((0,0)),D_{1}((0,0)))$, but not $D_{1}((0,1))$ for each unit (see Table 2). If, for instance, the observed treatment choices of unit $i$ in an exposed group are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,1))=1$, and the counterfactual treatment choice is $D_{i,1}((0,0))=0$, that unit belongs to the type $CM^{Z}\land NT^{T}$: unit $i$ is induced to the treatment by instrument but not by time. If the observed treatment choices of unit $i$ in an unexposed group are $D_{i,0}((0,0))=0$ and $D_{i,1}((0,0))=1$, and the counterfactual treatment choice is $D_{i,1}((0,1))=1$, that unit belongs to the type $AT^{Z}\land CM^{T}$: unit $i$ is induced to the treatment by time but not by instrument. Here, Assumption 3 (monotonicity assumption) excludes the defiers $DF^{Z}$ in exposed and unexposed groups respectively, written in gray color in both tables.
We now consider the interpretation of the parallel trends assumption in the outcome. Let $\Delta_{0}$ and $\Delta_{1}$ denote the average time trends of the unexposed outcome in exposed and unexposed groups, respectively.
$\Delta_{0}\equiv E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=0],$
$\Delta_{1}\equiv E[Y_{i,1}(D_{i,1}((0,0)))-Y_{i,0}(D_{i,0}((0,0)))|E_{i}=1].$
Note that $\Delta_{0}$ is estimable by the observed change in the outcome in an unexposed group during the two periods. By contrast, $\Delta_{1}$ is fundamentally unobservable because exposed group is assigned to the instrument path $z=(0,1)$.
To make the interpretations of $\Delta_{0}$ and $\Delta_{1}$ clear, we introduce the additional notation. For each time group $g^{t}\in\{AT^{T},NT^{T},CM^{T},DF^{T}\}$, let $w_{g^{t},e}\equiv Pr[G_{i}^{T}=g^{t}|E_{i}=e]$ denote the population share in group $E_{i}=e\in\{0,1\}$. Let $\Delta_{g^{t},e}$ denote the expectation of the potential outcome path in the absence of a policy shock for time group $g^{t}$ in group $E_{i}=e$:
$\Delta_{AT^{T},e}\equiv E[Y_{i,1}(1)-Y_{i,0}(1)|AT^{T},E_{i}=e],$ $\Delta_{NT^{T},e}\equiv E[Y_{i,1}(0)-Y_{i,0}(0)|NT^{T},E_{i}=e],$
$\Delta_{CM^{T},e}\equiv E[Y_{i,1}(1)-Y_{i,0}(0)|CM^{T},E_{i}=e],$ $\Delta_{DF^{T},e}\equiv E[Y_{i,1}(0)-Y_{i,0}(1)|DF^{T},E_{i}=e].$
We can interpret each $\Delta_{g^{t},e}$ as the effect of time on the outcome through treatment for time group $g^{t}$ in group $E_{i}=e$. First, $\Delta_{AT^{T},e}$ and $\Delta_{NT^{T},e}$ only reflect the expectation of the time effect for treated and untreated potential outcomes, respectively.
$\Delta_{AT^{T},e}$ $=\underbrace{E[Y_{i,1}(1)-Y_{i,0}(1)|AT^{T},E_{i}=e]}_{\text{Timeeffect among $AT^{T}$}},\hskip 5.69054pt\Delta_{NT^{T},e}$ $=\underbrace{E[Y_{i,1}(0)-Y_{i,1}(0)|NT^{T},E_{i}=e]}_{\text{Timeeffect among $NT^{T}$}}.$
This is because the time always-takers $AT^{T}$ and the time never-takers $NT^{T}$ do not change their treatment status during the two periods. Next, $\Delta_{CM^{T},e}$ and $\Delta_{DF^{T},e}$ can be decomposed into two terms.
$\Delta_{CM^{T},e}$ $=E[Y_{i,1}(0)-Y_{i,0}(0)|CM^{T},E_{i}=e]+\underbrace{E[Y_{i,1}(1)-Y_{i,1}(0)|CM^{T},E_{i}=e]}_{\text{Selection gain among $CM^{T}$}},$
$\Delta_{DF^{T},e}$ $=E[Y_{i,1}(1)-Y_{i,0}(1)|DF^{T},E_{i}=e]+\underbrace{E[Y_{i,1}(0)-Y_{i,1}(1)|DF^{T},E_{i}=e]}_{\text{Selection gain among $DF^{T}$}}.$
The first terms in $\Delta_{CM^{T},0}$ and $\Delta_{DF^{T},0}$ represent the average time effect for untreated and treated potential outcomes in each time group in group $E_{i}=e$, respectively. The second term represents the selection gain in each time group in group $E_{i}=e$, respectively. The latter terms arise because the time compliers $CM^{T}$ and the time defiers $DF^{T}$ are affected by time and change their treatment status from period 00 to period $1$.
The following theorem establishes the relationship between $\Delta_{e}$ and $\Delta_{g^{t},e}$.
Theorem 2.
Suppose Assumptions 1-4 hold, then
$\Delta_{e}=\sum_{g^{t}}w_{g^{t},e}\Delta_{g^{t},e},$
where the weight $w_{g^{t},e}$ is non-negative and sums to one; that is, $\sum_{g^{t}}w_{g^{t},e}=1$ holds.
Proof.
This directly follows from the Law of Iterated Expectation:
$\Delta_{e}$ $=E[Y_{1}(1)-Y_{0}(1)|AT^{T},E=e]Pr(AT^{T}|E=e)$
$+E[Y_{1}(0)-Y_{0}(0)|NT^{T},E=e]Pr(NT^{T}|E=e)$
$+E[Y_{1}(1)-Y_{0}(0)|CM^{T},E=e]Pr(CM^{T}|E=e)$
$+E[Y_{1}(0)-Y_{0}(1)|DF^{T},E=e]Pr(DF^{T}|E=e)$
$=\sum_{g^{t}}w_{g^{t},e}\Delta_{g^{t},e}.$
Clearly, the weight $w_{g^{t},e}$ sums to one by construction. ∎
Theorem 2 shows that the average time trend of the unexposed outcome in group $e$ can be written as the weighted average of the expectation of the potential outcomes, and can be interpreted as measuring the convex average of the time effect for each time group $g^{t}$ in group $e$. The weight assigned to each $\Delta_{g^{t},e}$ is natural in that it reflects the population share in each time group $g^{t}$ in group $e$. Hereafter, we call $\Delta_{e}$ an expected time gain in group $e$.
From Theorem 2, we can interpret that the parallel trends assumption in the outcome (Assumption 6), $\Delta_{0}=\Delta_{1}$, requires the same expected time gain between exposed and unexposed groups; that is, the effects of time on the outcome through treatment are the same between the two groups.
Assumption 6 is required for DID-IV identification strategy because in the numerator of the Wald-DID estimand, we should isolate the effect of the instrument on the outcome through treatment from that of time in an exposed group. The underlying idea to capture the former effect is parallel to that in DID designs: we purge the latter effect by exploiting the observed time trend of the outcome in the group that did not receive the policy shock (instrument) in the second period.
Finally, we provide an empirical example to illustrate the interpretation of each notation and Assumption 6 in practice.
Example.
Suppose that the treatment $D_{i,t}\in\{0,1\}$ represents participation in a job training program, the outcome $Y_{i,t}$ represents real earnings, and the instrument $Z_{i,1}\in\{0,1\}$ represents a policy shock that encourages people to receive the treatment in the second period. [^10] Assume that the treatment is non-staggered and the treatment and the instrument are not randomized across units. In this setting, the instrumental variable strategy nor the canonical DID designs are directly applicable. In this setting, the time compliers $CM^{T}$ are the units who start adopting the program in the second period in the absence of the policy shock, and $w_{CM^{T},0}$ measures its share in an unexposed group. $\Delta_{CM^{T},0}$ represents the average time trends of real earnings among the time compliers in an unexposed group, and can be decomposed into the time effects on real earnings without policy shock, and the causal effects of program participation on real earnings in the second period, as their treatment adoption behavior is affected by time. Assumption 6 requires that the average time trends of real earnings are the same without policy shock between exposed and unexposed groups. This assumption can be interpreted as the restriction on the effect of time on real earnings through program participation between the two groups.
2.5   Comparing DID-IV to Fuzzy DID
de Chaisemartin and D’Haultfœuille (2018) also investigate the identification assumptions for the Wald-DID estimand to capture the causal effects under the same setting as in this article. [^11] The Wald-DID estimand considered in dCDH corresponds to ours in this article. In their empirical application, dCDH revisit Duflo (2001), in which the group variable $E_{i}$ ($G_{i}$ in dCDH) takes one if unit $i$ belongs to the region that are highly exposed to the policy shock. However, they formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs. In this section, we compare DID-IV to Fuzzy DID and point out the issues embedded in Fuzzy DID in an overview of the detailed discussion provided in online Appendix Section E.

### v6

@@ 2  DID-IV in two time periods
In this section, we formalize an instrumented difference-in-differences (DID-IV) in two-period/two-group settings. Specifically, we establish the target parameter and the identifying assumptions in this design. At the end of this section, we also discuss the connections between DID-IV and Fuzzy DID proposed by de Chaisemartin and D’Haultfœuille (2018).
@@ 2.1  Set up
In practice, researchers are interested in the effect of a treatment $D_{i,t}$ on an outcome $Y_{i,t}$, and the instrument $Z_{i,1}$ typically represents a policy shock that encourages people to adopt the treatment in period $t=1$. For instance, Duflo (2001) estimates returns to schooling in Indonesia, exploiting variation arising from a new school construction program across regions and cohorts as an instrument for education attainment.
Next, we introduce the potential outcomes framework. Let $Y_{i,t}(d,z)$ denote the potential outcome in period $t$ when unit $i$ receives the treatment path $d\in\mathcal{S}(D)$ and the instrument path $z\in\mathcal{S}(Z)$. Similarly, let $D_{i,t}(z)$ denote the potential treatment status in period $t$ when unit $i$ receives the instrument path $z\in\mathcal{S}(Z)$. We refer to $D_{i,t}((0,0))$ as unexposed treatment and $D_{i,t}((0,1))$ as exposed treatment. Since the treatment and the instrument take only two values, one can write the observed treatments $D_{i,t}$ and outcomes $Y_{i,t}$ as follows.
@@
Assumption 1 states that potential outcomes $Y_{i,t}(d,z)$ depend only on the current treatment status $d_{t}$ and the instrument path $z$. In the DID literature, several studies adopt this assumption in non-staggered treatment settings (e.g. de Chaisemartin and D’Haultfœuille (2020); Imai and Kim (2021)).
@@ 2.2  Target parameter in $2\times 2$ DID-IV designs
This parameter measures the treatment effects in period $1$ for units that belong to the exposed group ($E_{i}=1$) and are induced to receive the treatment by the instrument in that period. In the DID-IV literature, de Chaisemartin (2010) considers the same target parameter, while Hudson et al. (2017) define the local average treatment effect (LATE) in period t=1t=1—unconditional on $E_{i}$—as their target parameter. The LATET has been also proposed in heterogeneous effects IV models with binary instrument (e.g. Słoczyński (2020); Słoczyński et al. (2022)).
We focus on the LATET for two reasons. First, this parameter would be particularly of interest if the instrument reflects a policy change of interest to the researcher (Heckman and Vytlacil (2001), J. and Vytlacil,Edward (2005)). Second, the LATET is a natural extension of the target parameter in DID designs, the so-called average treatment effects on the treated (ATT); both causal parameters measure the treatment effects among the units affected by a policy shock (the treatment in DID designs) and belonging to an exposed group (treatment group).
@@ 2.3  Identification assumptions in $2\times 2$ DID-IV designs
Given Assumptions 1-2, the observed outcome $Y_{i,t}$ can be written as
@@
Assumptions 1-2 also allow us to introduce the notions of exposed and unexposed outcomes. For any $z\in\mathcal{S}(Z)$, let $Y_{i,t}(D_{i,t}(z))$ denote the outcome if the instrument path is $z$:
@@
Hereafter, we refer to $Y_{i,t}(D_{i,t}((0,0)))$ as the unexposed outcomes and $Y_{i,t}(D_{i,t}((0,1)))$ as the exposed outcomes. [^6] Exposed and unexposed outcomes are not new concepts in econometrics. For example, in IV designs, the numerator of the Wald estimand compares expected outcomes across instrument values: E[Y|Z=1]E[Y|Z=0]E[Y|Z=1]-E[Y|Z=0]. This difference can be written as E[Y(D(1))|Z=1]E[Y(D(0))|Z=0]E[Y(D(1))|Z=1]-E[Y(D(0))|Z=0], which corresponds to a comparison between exposed and unexposed outcomes.
@@
Assumption 6 is a parallel trends assumption in the outcome, requiring that the evolution of the unexposed outcome is, on average, the same between exposed and unexposed groups.999de Chaisemartin et al. (2024) note that this assumption imposes restrictions on treatment effect heterogeneity when the standard parallel trends assumption holds between exposed and unexposed groups. For instance, in Duflo (2001), this assumption requires that the expectation of log annual earnings would have followed the same path from period 0 to period $1$ between exposed and unexposed groups in the absence of a policy shock.
When empirical researchers exploit variation arising from a policy shock as an instrument for treatment and use the Wald-DID estimand, they often refer to Assumptions 5 and 6. For instance, Duflo (2001) estimates returns to schooling in Indonesia, relying on “the identification assumption that the evolution of wages and education across cohorts would not have varied systematically from one region to another in the absence of the program”.
Finally, we assume a relevance condition. This assumption guarantees that the Wald-DID estimand $w_{DID}$ is well defined.
@@
2.4  Comparing DID-IV with Fuzzy DID
de Chaisemartin and D’Haultfœuille (2018) (henceforth, “dCDH”) also investigate the identifying assumptions for the Wald-DID estimand to capture the causal effects under the same setting considered here.101010dCDH also consider the DID-IV setting as in this paper. This follows from two observations. First, the group variable $G$ is included in their treatment participation equation $D=\mathbf{1}\{V\geq v_{GT}\}$ (see Assumption 3 in dCDH), implying that $G$ plays the role of the instrument $Z$ in our framework. Second, dCDH also assume the sharp assignment of the instrument: in their treatment participation equation, they impose $v_{10}=v_{00}$ (see page $1019$ of dCDH). However, dCDH formalize $2\times 2$ DID-IV designs differently, calling them Fuzzy DID designs. Moreover, dCDH point out that the Wald-DID estimand requires the stable treatment effect assumption to identify their target parameter.
In this section, we clarify the differences between this paper and dCDH, and provide the implications of these differences for treatment adoption behavior, the interpretation of the target parameter, and the use of the Wald-DID estimand. The detailed discussions and proofs are provided in Online Appendix D.
Since dCDH formalize Fuzzy DID designs under repeated cross section data, we first introduce notation for the repeated cross section setting. Let $Y$ and $D$ denote the outcome and the treatment, respectively. Let $Z\in\{0,1\}$ denote the group indicator, where $Z=1$ indicates the exposed group. Let $T\in\{0,1\}$ denote the time indicator. Let $Y(0),Y(1)$ and $D(0),D(1)$ denote the potential outcomes and the potential treatment choices, respectively. [^11] Here, we impose the no carryover assumption and the exclusion restriction on potential outcomes, and the no anticipation assumption on potential treatment choices, as in dCDH. Let $D_{t}(0)$ and $D_{t}(1)$ denote the potential treatment choices in period $T=t$ when the group indicator takes value $Z=0$ and $Z=1$, respectively.121212dCDH introduce the treatment participation equation $D=\mathbf{1}\{V\geq v_{ZT}\}$ and define the potential treatment choice in period $t$ as $D(t)=\mathbf{1}\{V\geq v_{Zt}\}$. In this paper, we instead adopt the notation $D_{t}(Z)=\mathbf{1}\{V\geq v_{Zt}\}$ to facilitate interpretation. In dCDH, the group variable $G$ plays the role of the instrument $Z$ in this paper. Accordingly, we relabel $G$ as $Z$ in the following discussion.